# Efficient multi-prompt evaluation of LLMs

### v1

@@ Abstract
Most popular benchmarks for comparing LLMs rely on a limited set of prompt templates, which may not fully capture the LLMs’ abilities and can affect the reproducibility of results on leaderboards. Many recent works empirically verify prompt sensitivity and advocate for changes in LLM evaluation. In this paper, we consider the problem of estimating the performance distribution across many prompt variants instead of finding a single prompt to evaluate with. We introduce PromptEval, a method for estimating performance across a large set of prompts borrowing strength across prompts and examples to produce accurate estimates under practical evaluation budgets. The resulting distribution can be used to obtain performance quantiles to construct various robust performance metrics (e.g., top 95% quantile or median). We prove that PromptEval consistently estimates the performance distribution and demonstrate its efficacy empirically on three prominent LLM benchmarks: MMLU, BIG-bench Hard, and LMentry. For example, PromptEval can accurately estimate performance quantiles across 100 prompt templates on MMLU with a budget equivalent to two single-prompt evaluations [^1] Our code and data can be found in https://github.com/felipemaiapolo/prompt-eval..
@@ 1.1 Related work
LLMs’ sensitivity to prompt templates. The sensitivity of Large Language Models (LLMs) to the prompts is well-documented. For example, Sclar et al. (2023) revealed that subtle variations in prompt templates in few-shot settings can lead to significant performance discrepancies among several open-source LLMs, with differences as large as 76 accuracy points in tasks from the SuperNaturalInstructions dataset (Wang et al., 2022). Additionally, they report that the performance of different prompt templates tends to correlate weakly between models. This finding challenges the reliability of evaluation methods that depend on a single prompt template. To measure LLMs sensitivity, the researchers suggested calculating a “performance spread,” which represents the difference between the best and worst performances observed. Mizrahi et al. (2023) conducted a complementary analysis using state-of-the-art models and subsets of BigBench and LMentry (Srivastava et al., 2022; Efrat et al., 2022). The authors arrive at similar conclusions with respect to LLMs’ sensitivity to the used prompt templates and empirically showed that the LLM ranking considering different formats are usually weakly or intermediately correlated with each other. As a solution to the lack of robustness in LLM evaluation, the authors propose the use of summary statistics, as the average performance, for LLM evaluation. Some other works, e.g., Voronov et al. (2024); Weber et al. (2023b, a), show that even when in-context examples are given to the models, the prompt templates can have a big impact on the final numbers, sometimes reducing the performance of the strongest model in their analyses to a random guess level (Voronov et al., 2024). In a different direction, Shi et al. (2024) acknowledges that different prompt templates have different performances and proposes using best-arm-identification to efficiently select the best template for an application at hand. One major bottleneck is still on how to efficiently compute the performance distribution for LLMs over many prompt templates; we tackle this problem.
@@
Efficient evaluation of LLMs. The escalating size of models and datasets has led to increased evaluation costs. To streamline evaluations, Ye et al. (2023b) considered minimizing the number of tasks within Big-bench (Srivastava et al., 2022). Additionally, Perlitz et al. (2023) observed that evaluations on HELM (Liang et al., 2022) rely on diversity across datasets, though the quantity of examples currently utilized is unnecessarily large. Perlitz et al. (2023) also highlighted the problems in evaluating with insufficient prompts and called to evaluate on more, suggesting evaluating the typical behavior by sampling prompts and examples together. To accelerate evaluations for classification tasks, Vivek et al. (2023) suggested clustering evaluation examples based on model confidence in the correct class. More recently, Polo et al. (2024) empirically showed that it is possible to shrink the size of modern LLM benchmarks and still retain good estimates for LLMs’ performances. Despite these advancements in streamlining LLM evaluations, there are no other works that propose a general and efficient multi-prompt evaluation method to the best of our knowledge.
@@
Item response theory (IRT). IRT (Cai et al., 2016; Van der Linden, 2018; Brzezińska, 2020; Lord et al., 1968) is a collection of statistical models initially developed in psychometrics to assess individuals’ latent abilities through standardized tests but with increasing importance in the fields of artificial intelligence and natural language processing (NLP). For example, Lalor et al. (2016) used IRT’s latent variables to measure language model abilities, Vania et al. (2021) applied IRT to benchmark language models and examine the saturation of benchmarks, and Rodriguez et al. (2021) explored various uses of IRT with language models, including predicting responses to unseen items, categorizing items by difficulty, and ranking models. Recently, Polo et al. (2024) suggested using IRT for efficient LLM performance evaluation, introducing the Performance-IRT (pIRT) estimator to evaluate LLMs. Our quantile estimation methodology is built upon pIRT.
@@ 2 Problem statement
In this section, we describe the setup we work on and what our objectives are. Consider that we want to evaluate a large language model (LLM) in a certain dataset composed of $J$ examples (questions/items) and each one of the examples is responded to by the LLM through prompting; we assume that there exists $I$ different prompt templates that can be used to evaluate the LLM. After the prompt template $i\in\mathcal{I}\triangleq[I]$ and example $j\in\mathcal{J}\triangleq[J]$ are channelled through the LLM, some grading system generates a correctness score $Y_{ij}\in\{0,1\}$, which assumes $1$ when the prompt template $i$ and example $j$ jointly yield a correct response and 00 otherwise [^2] In some cases, the correctness score may be a bounded number instead of binary – see Appendix B.. For each one of the prompt templates $i\in\mathcal{I}$, we can define its performance score as
@@ 5 Assessing multi-prompt evaluation strategies
We consider different variations of the model presented in (3.2) coupled with Algorithm 1; for all variations, we use linear $f_{\psi}$ and $g_{\gamma}$. The most basic version of the model in (3.2) assumes $x_{i}$ and $z_{j}$ are one-hot encoded vectors, i.e., vector of zeros with 1’s on the entries $i$ and $j$, reverting the model to a Rasch model (Georg, 1960; Chen et al., 2023). Despite its simplicity, we show that it can perform well in some cases. A more advanced instance of (3.2) assumes $x_{i}$ are either obtained using a sentence transformer (Reimers and Gurevych, 2019) to embed prompt templates or by extracting discrete covariates from the text, e.g., as the presence of line breaks, colons etc.(see Appendix Table 2). An example of a prompt template for LMentry used by Mizrahi et al. (2023) is “Can {category} be used to classify all the {words} provided? Respond with either "yes" or "no".” Our method also allows using example covariates $z_{j}$, however, upon preliminary tests with sentence transformer we didn’t observe improvements and chose to use one-hot-encoded vectors as in the basic Rasch model to represent examples. Next we detail the methods for obtaining the prompt covariates:
@@
  • Distribution estimation. Our results for quantile estimation can be seen in Figure 2. We see that, in general, all variations of PromptEval, including its simplest version (PE-Rasch), can do much better in distribution estimation when compared to the baseline. Among our methods, the ones that use covariates are the best ones.

### v2

@@ Abstract
Most popular benchmarks for comparing LLMs rely on a limited set of prompt templates, which may not fully capture the LLMs’ abilities and can affect the reproducibility of results on leaderboards. Many recent works empirically verify prompt sensitivity and advocate for changes in LLM evaluation. In this paper, we consider the problem of estimating the performance distribution across many prompt variants instead of finding a single prompt to evaluate with. We introduce PromptEval, a method for estimating performance across a large set of prompts borrowing strength across prompts and examples to produce accurate estimates under practical evaluation budgets. The resulting distribution can be used to obtain performance quantiles to construct various robust performance metrics (e.g., top 95% quantile or median). We prove that PromptEval consistently estimates the performance distribution and demonstrate its efficacy empirically on three prominent LLM benchmarks: MMLU, BIG-bench Hard, and LMentry. For example, PromptEval can accurately estimate performance quantiles across 100 prompt templates on MMLU with a budget equivalent to two single-prompt evaluations [^1] Our code can be found in https://github.com/felipemaiapolo/prompteval. The MMLU data can be found in https://huggingface.co/PromptEval..
@@ 1.1 Related work
LLMs’ sensitivity to prompt templates The sensitivity of Large Language Models (LLMs) to the prompts is well-documented. For example, Sclar et al. (2023) revealed that subtle variations in prompt templates in few-shot settings can lead to significant performance discrepancies among several open-source LLMs, with differences as large as 76 accuracy points in tasks from the SuperNaturalInstructions dataset (Wang et al., 2022). Additionally, they report that the performance of different prompt templates tends to correlate weakly between models. This finding challenges the reliability of evaluation methods that depend on a single prompt template. To measure LLMs sensitivity, the researchers suggested calculating a “performance spread,” which represents the difference between the best and worst performances observed. Mizrahi et al. (2023) conducted a complementary analysis using state-of-the-art models and subsets of BigBench and LMentry (Srivastava et al., 2022; Efrat et al., 2022). The authors arrive at similar conclusions with respect to LLMs’ sensitivity to the used prompt templates and empirically showed that the LLM ranking considering different formats are usually weakly or intermediately correlated with each other. As a solution to the lack of robustness in LLM evaluation, the authors propose the use of summary statistics, as the average performance, for LLM evaluation. Some other works, e.g., Voronov et al. (2024); Weber et al. (2023b, a), show that even when in-context examples are given to the models, the prompt templates can have a big impact on the final numbers, sometimes reducing the performance of the strongest model in their analyses to a random guess level (Voronov et al., 2024). In a different direction, Shi et al. (2024) acknowledges that different prompt templates have different performances and proposes using best-arm-identification to efficiently select the best template for an application at hand. One major bottleneck is still on how to efficiently compute the performance distribution for LLMs over many prompt templates; we tackle this problem.
@@
Efficient evaluation of LLMs The escalating size of models and datasets has led to increased evaluation costs. To streamline evaluations, Ye et al. (2023b) considered minimizing the number of tasks within Big-bench (Srivastava et al., 2022). Additionally, Perlitz et al. (2023) observed that evaluations on HELM (Liang et al., 2022) rely on diversity across datasets, though the quantity of examples currently utilized is unnecessarily large. Perlitz et al. (2023) also highlighted the problems in evaluating with insufficient prompts and called to evaluate on more, suggesting evaluating the typical behavior by sampling prompts and examples together by employing stratified sampling, where subscenarios give the strata; in our work, we also apply stratification but consider prompt templates and examples to give the strata. To accelerate evaluations for classification tasks, Vivek et al. (2023) suggested clustering evaluation examples based on model confidence in the correct class. More recently, Polo et al. (2024) empirically showed that it is possible to shrink the size of modern LLM benchmarks and still retain good estimates for LLMs’ performances. Similarly (and in parallel to this work) Ashury-Tahan et al. (2024) recognized unlabeled examples that better distinguish between models or prompts, by analyzing model outputs on them, hence saving costly annotation for them. Despite these advancements in streamlining LLM evaluations, there are no other works that propose a general and efficient method to estimate the benchmark performance of LLMs across prompt templates to the best of our knowledge.
@@
Item response theory (IRT) IRT (Cai et al., 2016; Van der Linden, 2018; Brzezińska, 2020; Lord et al., 1968) is a collection of statistical models initially developed in psychometrics to assess individuals’ latent abilities through standardized tests but with increasing importance in the fields of artificial intelligence and natural language processing (NLP). For example, Lalor et al. (2016) used IRT’s latent variables to measure language model abilities, Vania et al. (2021) applied IRT to benchmark language models and examine the saturation of benchmarks, and Rodriguez et al. (2021) explored various uses of IRT with language models, including predicting responses to unseen items, categorizing items by difficulty, and ranking models. Recently, Polo et al. (2024) suggested using IRT for efficient LLM performance evaluation, introducing the Performance-IRT (pIRT) estimator to evaluate LLMs. Our quantile estimation methodology is built upon pIRT.
@@ 2 Problem statement
In this section, we describe the setup we work on and what our objectives are. Consider that we want to evaluate a large language model (LLM) in a certain dataset composed of $J$ examples (also known as questions or items in the literature) and each one of the examples is responded to by the LLM through prompting; we assume that there exists $I$ different prompt templates that can be used to evaluate the LLM. After the prompt template $i\in\mathcal{I}\triangleq[I]$ and example $j\in\mathcal{J}\triangleq[J]$ are channelled through the LLM, some grading system generates a correctness score $Y_{ij}\in\{0,1\}$, which assumes $1$ when the prompt template $i$ and example $j$ jointly yield a correct response and 00 otherwise [^2] In some cases, the correctness score may be a bounded number instead of binary – see Appendix B.. For each one of the prompt templates $i\in\mathcal{I}$, we can define its performance score as
@@ 5 Assessing multi-prompt evaluation strategies
We consider different variations of the model presented in (3.2) coupled with Algorithm 1; for all variations, we use linear $f_{\psi}$ and $g_{\gamma}$. The most basic version of the model in (3.2) assumes $x_{i}$ and $z_{j}$ are one-hot encoded vectors, i.e., vector of zeros with 1’s on the entries $i$ and $j$, reverting the model to a Rasch model (Georg, 1960; Chen et al., 2023). Despite its simplicity, we show that it can perform well in some cases. A more advanced instance of (3.2) assumes $x_{i}$ are either obtained using a sentence transformer (Reimers and Gurevych, 2019) to embed prompt templates or by extracting discrete covariates from the text, e.g., as the presence of line breaks, colons etc.(see Appendix Table 2). An example of a prompt template for LMentry used by Mizrahi et al. (2023) is “Can {category} be used to classify all the {words} provided? Respond with either "yes" or "no".” Our method also allows using example covariates $z_{j}$, however, upon preliminary tests with sentence transformer we didn’t observe improvements and chose to use one-hot-encoded vectors as in the basic Rasch model to represent examples. Next, we detail the methods for obtaining the prompt covariates:
@@
  • Distribution estimation. Our results for distribution estimation can be seen in Figure 2. We see that, in general, all variations of PromptEval, including its simplest version (PE-Rasch), can do much better in distribution estimation when compared to the baseline. Among our methods, the ones that use covariates are the best ones.
@@ Appendix G Details MMLU data
Next, we utilize the unitxt  [Bandel et al., 2024] preprocessing library to build custom datasets with the generated templates. Standardized and accurate evaluation is then carried out via the LM-Eval-Harness  [Gao et al., 2023] evaluation library.